# GenCode: A Generic Data Augmentation Framework for Boosting Deep Learning-Based Code Understanding

### v2

@@ Introduction
2:$\mathcal{P},Y$: original code data and labels
3:$\mathcal{T}=\left\{t\right\}:$ a set of code augmentation methods
@@
6:     $\mathcal{P^{\prime}},Y^{\prime}\ \leftarrow\phi,\phi$triangleright initialization
@@
9:         $Y^{\prime}\leftarrow Y^{\prime}\cup Y$triangleright code augmentation
10:     $I\leftarrow\mathsf{Importance}\left(M,\mathcal{P^{\prime}}\right)$triangleright computing importance
11:     $Index\leftarrow\mathsf{argSort}\left(I\right)\left[:||\mathcal{P}||\right]$triangleright sort by importance
12:     $X_{train}\leftarrow\mathcal{P^{\prime}}\left[Indext\right]$
13:     $Y_{train}\leftarrow Y^{\prime}\left[Indext\right]$
@@
Search Space
Even though the code has strict syntax constraints, recent work (Dong et al. 2023b) found that these data augmentation methods that slightly break the syntax of the source code e.g., Random Swap, can still be useful for code model training. To enhance the diversity of source code within the feature space, therefore, GenCode considers two types of code augmentation methods, semantic/syntax-preserving methods, and syntax-breaking methods (see Definition 4 and Definition 5).
GenCode adapts code refactoring methods as semantic/syntax-preserving methods. Code refactoring methods follow the constraints of programming languages to edit the code while keeping the main functionality of the code. On the other hand, text augmentation methods are used as syntax-breaking methods in GenCode which are widely used in the NLP field. For example, text augmentation methods can randomly delete some tokens in the code. However, even though the syntax has been broken, the code model can still successfully predict those codes since it reads the code as a sequence. In total, as shown in Table 2, GenCode currently supports 18 code refactoring methods and 5 text augmentation methods. Besides, Figure 1 provides an example of the training candidate that is to generate the search space. In the syntax-preserving, we present a code refactoring method named API Renaming, which involves modifying the function name from ”main()” to ”even()”. In the syntax-breaking, we demonstrate another instance of API Renaming, where two different statements within a program are randomly selected and their positions swapped. As depicted in this example, the third and fourth lines have been interchanged.
@@
Filed enhancement
Synonym replacement (SR) Random insertion (RI) Random swap (RS Random deletion (RD) Back-translation (BT)
@@
Table 3: Effectiveness of data augmentation methods w.r.t. test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data and ASR $\downarrow$ (%) on test data. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
GCJ Refactory Java250 BigCloneBench
Model DA method Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$
No Aug 90.96 ± 0.07 33.31 56.13 96.21 ± 0.13 36.31 42.31 96.36 ± 0.09 40.55 53.37 96.91 ± 0.12 9.74 24.61
Refactor 91.74 ± 0.11 35.56 52.84 95.95 ± 0.19 32.64 40.89 96.44 ± 0.12 38.35 52.56 96.97 ± 0.21 8.55 18.29
SR 70.66 ± 0.11 35.46 52.71 96.71 ± 0.11 39.25 53.91 96.21 ± 0.02 42.01 53.79 97.02 ± 0.15 12.52 26.11
RI 59.91 ± 0.21 40.78 65.76 96.43 ± 0.18 38.45 45.57 96.15 ± 0.06 43.15 55.07 97.38 ± 0.23 8.31 30.08
RS 93.24 ± 0.13 28.26 52.44 96.73 ± 0.15 38.63 45.98 96.49 ± 0.04 39.36 50.79 96.92 ± 0.12 13.41 30.13
RD 76.98 ± 0.09 35.65 55.46 96.46 ± 0.18 30.23 41.32 96.61 ± 0.12 38.31 52.42 97.11 ± 0.19 9.58 30.15
BT 82.74 ± 0.15 40.38 55.95 94.31 ± 0.38 62.98 69.87 96.06 ± 0.06 42.21 53.32 97.13 ± 0.17 11.29 26.32
MixCode 92.21 ± 0.25 29.28 52.53 96.76 ± 0.31 31.04 41.01 96.62 ± 0.23 36.14 50.67 - - -
CodeBERT GenCode 96.23 ± 0.18 22.33 46.62 99.32 ± 0.14 23.75 31.25 98.89 ± 0.09 29.89 40.65 98.93 ± 0.11 7.12 15.83
No Aug 93.96 ± 0.15 31.23 56.02 96.84 ± 0.16 22.97 30.26 96.41 ± 0.11 23.16 42.66 96.88 ± 0.09 4.19 12.22
Refactor 91.72 ± 0.21 30.31 56.58 95.49 ± 0.21 24.26 32.57 96.53 ± 0.15 23.13 42.42 97.11 ± 0.23 6.05 6.17
SR 89.46 ± 0.13 40.02 66.68 95.57 ± 0.13 28.39 47.15 96.46 ± 0.14 22.92 42.18 97.14 ± 0.16 6.13 10.16
RI 80.47 ± 0.21 43.51 65.75 94.51 ± 0.25 25.08 32.87 96.52 ± 0.09 22.37 42.33 97.35 ± 0.17 2.23 6.28
RS 94.76 ± 0.15 30.18 54.75 97.88 ± 0.15 21.85 29.67 96.44 ± 0.14 22.68 41.98 97.54 ± 0.12 6.03 11.94
RD 91.71 ± 0.26 31.73 60.92 96.53 ± 0.19 20.35 26.98 96.56 ± 0.21 22.39 41.84 97.16 ± 0.21 6.18 8.27
BT 80.47 ± 0.05 41.65 62.99 95.99 ± 0.27 35.56 46.23 96.39 ± 0.13 22.91 42.32 97.11 ± 0.16 8.18 14.12
MixCode 94.19 ± 0.25 30.26 54.49 97.29 ± 0.15 14.89 26.53 96.89 ± 0.19 20.03 42.01 - - -
GraphCodeBERT GenCode 97.93 ± 0.14 26.43 48.87 99.45 ± 0.17 15.78 21.92 99.16 ± 0.13 18.76 33.81 99.23 ± 0.08 1.86 4.62
No Aug 95.02 ± 0.07 30.78 55.67 97.22 ± 0.17 22.45 29.69 96.98 ± 0.08 31.76 49.82 97.38 ± 0.13 10.32 22.97
Refactor 92.21 ± 0.14 30.12 55.93 96.43 ± 0.22 28.76 35.45 96.92 ± 0.11 30.66 47.12 97.24 ± 0.12 9.32 22.05
SR 90.14 ± 0.21 39.21 51.84 94.67 ± 0.14 28.24 36.26 96.53 ± 0.05 28.65 46.01 97.31 ± 0.08 15.87 25.98
RI 78.77 ± 0.05 41.02 63.46 95.02 ± 0.15 24.78 33.61 96.11 ± 0.09 27.14 46.76 97.46 ± 0.13 10.43 29.76
RS 95.66 ± 0.14 29.35 51.74 97.97 ± 0.29 21.32 27.76 96.34 ± 0.11 28.45 48.45 97.03 ± 0.15 14.33 30.43
RD 92.01 ± 0.11 31.04 56.85 97.52 ± 0.26 19.93 24.87 97.21 ± 0.07 26.09 45.24 97.43 ± 0.07 9.98 28.87
BT 81.45 ± 0.13 39.65 54.24 95.38 ± 0.11 33.62 43.67 96.45 ± 0.14 29.47 45.38 97.65 ± 0.14 10.43 25.34
MixCode 94.32 ± 0.16 29.82 42.56 97.32 ± 0.13 12.49 18.32 96.91 ± 0.11 28.73 48.87 - - -
CodeT5 GenCode 98.85 ± 0.11 23.57 44.76 99.64 ± 0.06 12.13 19.56 99.38 ± 0.11 21.63 39.76 99.34 ± 0.12 7.39 12.15
Importance-Guided Data Selection
@@
After constructing the training candidate, we then choose the most valuable code data that can benefit the code model training. Based on our Finding drawn from the preliminary study, GenCode chooses the simplest way to measure the importance of code data to the model – the loss value. That means, GenCode takes all the transformed code as input to the code model and records all the loss values first, then ranks all the code in descending order based on the loss values and selects the first $K$ code as the training data. Following the common setting in the data augmentation problem (Dong et al. 2023a), $K$ is set as the number of original training data. GenCode chooses loss value as the indicator since it can be generalized to any code model and task.
Evaluation
@@
We evaluate both the accuracy and robustness of trained code models. Besides, we check how does importance score affects the effectiveness of GenCode.
Datasets and Models
@@
Datasets and Models We evaluate GenCode on four code understanding tasks, bug detection (Hu et al. 2019), authorship attribution (Yang et al. 2022), code clone detection (Yang et al. 2022) (Note that, MixCode is not suitable for this task since its input is code pairs.), and problem classification (Puri et al. 2021). For each code task, we fine-tune three types of models based on pre-trained code models, CodeBERT, GraphCodebERT, and CodeT5. For the model training, we follow existing works (Hu et al. 2019; Wang et al. 2021; Puri et al. 2021; Yang et al. 2022; Dong et al. 2023a), to set all configurations. Specifically, we set the training epoch as 50 for all tasks and models. Regarding the optimizer, we utilize Adam (Kingma and Ba 2014) with a learning rate of $10^{-3}$ for all the models. Besides, we use early stopping with a patience of 20 to mitigate overfitting. To reduce the influence of randomness, we train each model five times and report the average results along with the standard deviation. For MixCode, we follow the recommendation of the original paper and set the Mixup ratio as $\alpha=0.1$. For the setting of adversarial attacks, we directly use the default configurations provided by the original papers  (Yang et al. 2022).
@@
Evaluation Metrics. In the evaluation, we assess the performance of trained models from two perspectives, namely, clean accuracy and model robustness. For the robustness evaluation, we check the attack success rate (ASR) metric of two well-established adversarial attack methods, namely, ALERT (Yang et al. 2022), which uses variable renaming while considering natural semantics, and MHM (Zhang et al. 2020), which also randomly replaces variable names but does not account for natural semantics. We do not consider the natural robustness since the existing natural robustness datasets can not fully support our studied tasks, i.e., no natural robustness datasets support authorship attribution, bug detection, and code clone detection.
@@
Baseline Methods. We consider model training without data augmentation (No Aug), Refactor (Yu, Wang, and Wang 2022), a family of text-oriented data augmentation methods (Dong et al. 2023b), and MixCode as baseline methods. Refactor, which keeps code syntax rules, is a widely used data augmentation framework specifically designed for code data. In contrast, text-oriented code data augmentation methods, which slightly break syntax rules, have recently proven more effective than Refactor. Additionally, MixCode is a more recent and high-performing code data augmentation method that significantly enhances code models.
Accuracy Evaluation
@@
Overall accuracy. Column Accuracy in Table 3 summarizes the final test accuracy of code models trained by using different code data augmentation methods including GenCode. Surprisingly, from the results, we can see that most of the existing code augmentation methods except RS and MixCode negatively affect the performance of code models. This finding is similar to the conclusion from recent code augmentation studies (Bielik and Vechev 2020; Yu, Wang, and Wang 2022; Dong et al. 2023b) which states code augmentation has limited advantage in boosting the training of code models and also indicates that data augmentation for code is a challenging problem.
@@
Then, compared to the existing data augmentation methods, the results demonstrate that GenCode consistently enhances the training and achieves the best results across all the considered datasets and models. Concretely, on average, GenCode generates models with 4.52%, 3.05%, 2.56%, and 2.11% higher accuracy than No Aug in GCJ, Refactory, Java250, and BigCloneBench, respectively. Besides, compared to MixCode, GenCode can achieve 4.1%, 2.34%, and 2.33% better results in GCJ, Refactory, and Java250 datasets, respectively.
Statistical analysis. We conducted statistical tests, including the T-Test and Wilcoxon signed-rank test, across all pre-trained models. The results show that GenCode significantly (with p-value $<$ 0.05) outperforms other methods including MixCode. In conclusion, GenCode is effective in training accurate code models.
@@
Figure 2: Convergence speed of CodeBERT using different code data augmentation methods in each task.
Convergence speed analysis. Figure 2 depicts the average test accuracy at each training epoch of CodeBERT using different data augmentation methods. The results demonstrated that GenCode consistently outperforms other baseline methods during the whole training process. We can conclude that the choice of the number of training epochs will not affect the advantages of our method over other methods.
@@
Considering Figure 3, we can see that the embedding produced by GenCode is clearly distributed compared to other methods. Besides, we computed the average distance between the data points from different classes, i.e., the average distance between red points and blue points in Figure 3. The results are shown in the following Table. The results confirm that GenCode can better distinguish the two classes. Interestingly, we found that MixCode has the smallest distance result. That means even though MixCode can train code models with good performance, the code models have low confidence in their predictions, thus, the data are near the decision boundaries. This phenomenon is similar to the findings in (Verma et al. 2019), where the representations obtained by interpolating hidden states have two desirable properties: the class-representations are flattened into a minimal amount of directions of variation, and all points in between these flat representations, most unobserved during training and off the data manifold, are assigned low-confidence predictions.

### v3

@@ 5.1 Overview
[…]
@@
6:  $\mathcal{P^{\prime}},Y^{\prime}\ \leftarrow\phi,\phi$ $\triangleright$ initialization
@@
9:   $Y^{\prime}\leftarrow Y^{\prime}\cup Y$ $\triangleright$ code augmentation
10:  $I\leftarrow\mathsf{Importance}\left(M,\mathcal{P^{\prime}}\right)$ $\triangleright$ computing importance
11:  $Index\leftarrow\mathsf{argSort}\left(I\right)\left[:||\mathcal{P}||\right]$ $\triangleright$ sort by importance
12:  $X_{train}\leftarrow\mathcal{P^{\prime}}\left[Index\right]$
13:  $Y_{train}\leftarrow Y^{\prime}\left[Index\right]$
@@
More specifically, Fig. 2 depicts the workflow of GenCode in one training epoch. First, a series of code augmentation methods are applied to the original dataset to generate an augmented dataset (search space). Then, samples are selected based on their influence score, i.e., the loss during model training (importance-guided data selection), ensuring the selected dataset keeps the same size as the original dataset for a fair comparison.
Figure 2: Workflow of GenCode in one training epoch.
Overall, GenCode contains two main steps, utilizing code augmentation methods to compose the search space and importance-guided data selection (will be introduced in Section 5.2 and Section 5.3, respectively).
5.2 Search Space
Even though the code has strict syntax constraints, recent work Dong et al. (2025) found that these data augmentation methods that slightly break the syntax of the source code e.g., Random Swap, can still be useful for code model training. To enhance the diversity of source code within the feature space, GenCode considers two types of code augmentation methods, semantic/syntax-preserving methods, and syntax-breaking methods (see Definition 4 and Definition 5). For instance, synonym replacement (SR) may turn an identifier ”count” into a natural-language word, such as ”amount”, which is not declared in the symbol table and thus results in a compilation error. In contrast, refactoring-based renaming updates all references consistently and guarantees syntactic correctness.
GenCode adapts code refactoring methods (i.e., API renaming and Duplication) as semantic/syntax-preserving methods. Code refactoring methods follow the constraints of programming languages (PL) to edit the code while keeping the main functionality of the code. On the other hand, text-oriented data augmentation methods Dong et al. (2025) are used as syntax-breaking methods in GenCode, which are widely used in the NLP field. For example, text-oriented data augmentation methods can randomly delete some tokens in the code. However, even though the syntax has been broken, the code model can still successfully predict those codes since it reads the code as a sequence. In total, as shown in Table 1, GenCode currently supports 18 code refactoring methods and 5 text-oriented data augmentation methods. Besides, Fig. 3 provides an example of the training candidate that is to generate the search space. In the syntax-preserving, we present a code refactoring method named API Renaming, which involves modifying the function name from ”main()” to ”even()”. In the syntax-breaking, we demonstrate another instance of Random Swap, where two different statements within a program are randomly selected and their positions swapped. As depicted in this example, the third and fourth lines have been interchanged.
@@ 5.2 Search Space
Field enhancement
Synonym replacement (SR) Random insertion (RI) Random swap (RS) Random deletion (RD) Back-translation (BT)
@@
5.3 Importance-Guided Data Selection
@@
After constructing the training candidate, we then choose the most valuable code data that can benefit the code model training. Based on our Finding drawn from the preliminary study, GenCode chooses the simplest way to measure the importance of code data to the model – the loss value. That means, GenCode takes all the transformed code as input to the code model and records all the loss values first, then ranks all the code in descending order based on the loss values and selects the first $K$ code as the training data. Following the common setting in the data augmentation problem Dong et al. (2023), $K$ is set as the number of original training data. GenCode chooses loss value as the indicator since it can be generalized to any code model and task.
6 Evaluation Setup
@@
We evaluate both the accuracy and robustness of trained code models. Besides, we check how does influence score affects the effectiveness of GenCode.
6.1 Study Design
We conduct experiments to evaluate the usefulness of GenCode and answer the following research questions.
RQ1: How effective is GenCode in producing accurate code models? Accuracy is the fundamental metric that is used to evaluate the performance of a trained model. Therefore, we first study the basic DNN property, the test accuracy, of each selected pre-trained code model. Specifically, we compare GenCode with models trained without using data augmentation and models trained by baseline data augmentation methods (will be introduced in Section 6.4), and we additionally provide the evaluation results of text-oriented data augmentation methods (as mentioned in Section 5.2) to demonstrate the usefulness of GenCode. Additionally, to ensure a fair comparison in terms of training cost, we apply each data augmentation method to every sample in the original training dataset, resulting in an augmented dataset of the same size. These augmented datasets are fixed prior to training and remain unchanged across all training epochs.
RQ2: How effective is GenCode in producing robust code models? Robustness is another important property of DNN models. In this study, we mainly focus on adversarial robustness and natural robustness. Specifically, for the adversarial robustness, we utilize two widely used adversarial attack methods including MHM Zhang et al. (2020) and ALERT Yang et al. (2022), to assess the performance of code models trained by the data augmentation method in robustness improvement. For natural robustness, we randomly apply a single code refactoring method Dong et al. (2023) to each sample in order to construct a transformed code dataset that preserves the original semantic labels, thereby enabling a comprehensive evaluation of the generalization capability under natural variations.
RQ3: How does the influence score affect the effectiveness of GenCode? The importance of score-based data selection is the key component is GenCode. Therefore, in this RQ, we change the selection strategies, i.e., selecting data with minimum confidence and randomly selecting data, to study the contribution of our proposed selection strategy. To achieve this, we repeat the experiments outlined in RQ1 and RQ2. Subsequently, we investigate how the selection of different scores influences the performance of pre-trained models. Note that, for a fair comparison, we evaluate only the adversarial robustness, as the use of random may artificially inflate performance on the transformed dataset Dong et al. (2023).
RQ4: How does the data selection method affect the effectiveness of GenCode? Data selection strategies play a critical role in determining the performance of trained code models Hu et al. (2024). In this RQ, we select representative methods from the literature to serve as baselines for comparison with GenCode. Specifically, we include LIME Ribeiro et al. (2016), which provides interpretable, local approximations of complex model predictions and can be leveraged during the model training phase to inform data selection by identifying influential samples. As one of the classical data selection methods, Influence Function Koh and Liang (2017) estimates the effect of individual training samples by approximating how infinitesimal perturbations to the training data impact the learned parameters, and can be effectively utilized during the phase of model training to guide data selection. Additionally, we incorporate BADGE Ash et al. (2019), which selects data by applying K-means clustering on gradient embeddings, and has been shown to be a stable and effective active learning method in recent empirical studiesHu et al. (2024). We evaluate the performance of selected data selection methods in terms of both accuracy and natural robustness. Adversarial robustness is not considered in this study, as existing adversarial attacks in RQ2 are not well-suited for LLMs.
6.2 Datasets and Models
@@
We evaluate GenCode on four code understanding tasks as shown in Table 2, bug detection Hu et al. (2019), authorship attribution Yang et al. (2022), code clone detection Yang et al. (2022) (Note that, MixCode is not suitable for this task since its input is code pairs.), and problem classification Puri et al. (2021). We conduct all experiments on a server equipped with 4 NVIDIA RTX A6000 GPUs. For each code task, we fine-tune three types of models based on pre-trained code models, CodeBERT, GraphCodeBERT, and CodeT5. For the model training, we follow existing works Hu et al. (2019); Wang et al. (2021); Puri et al. (2021); Yang et al. (2022); Dong et al. (2023), to set all configurations. Specifically, we set the training epoch as 50 for all tasks and models. Regarding the optimizer, we utilize Adam Kingma and Ba (2014) with a learning rate of $10^{-3}$ for all the models. Besides, we use early stopping with a patience of 20 to mitigate overfitting. To reduce the influence of randomness, we train each model five times and report the average results along with the standard deviation. For MixCode, we follow the recommendation of the original paper and set the Mixup ratio as $\alpha=0.1$. For the setting of adversarial attacks, we directly use the default configurations provided by the original papers  Yang et al. (2022). Considering the code-specific Large Language Models (LLMs), we select StarCoder2 Lozhkov et al. (2024), a representative model from the BigCode  [^6] https://www.bigcode-project.org project, which is a collaborative initiative dedicated to the responsible development of LLMs for Code, alongside the recently released Qwen2.5-Coder Hui et al. (2024).
Table 2: Details of tasks, datasets, and DNN models. #Training and #Test are the number of data for training and testing, respectively.
Dataset Language Task #Training #Test Model
Google Code Jam (GCJ) Python Authorship attribution 528 132 CodeBERT GraphCodeBERT CodeT5 StarCoder2-7B Qwen2.5-Coder-7B
Refactory Python Bug detection 3,380 423
Java250 Java Problem classification 48,000 15,000 CodeBERT GraphCodeBERT CodeT5 StarCoder2-7B Qwen2.5-Coder-7B
BigCloneBench Java Clone detection 90,102 4,000
6.3 Configuration
We conduct all experiments on a computation server equipped with 4× NVIDIA A100 (40GB) GPUs, two Intel Xeon Gold 6330 CPUs, and 512GB RAM. To ensure fairness and consistency across benchmarks, we adopt a unified training configuration following established practices in prior work Wang et al. (2021); Dong et al. (2023); Hu et al. (2019); Yang et al. (2022); Puri et al. (2021). Specifically, we train CodeBERT, GraphCodeBERT, and CodeT5 for 50 epochs, and train StarCoder2-7B and Qwen2.5-Coder-7B for 20 epochs, using the Adam optimizer (learning rate = 1e-3, weight decay = 0.01). The batch size is fixed to 32, and the maximum sequence length is set to 512 tokens for all encoder-based architectures. To mitigate overfitting and ensure stable convergence, we apply early stopping with a patience of 20 epochs. Each experiment is executed five independent times, and we report the mean and standard deviation to account for stochastic variability. For MixCode, we follow the settings recommended in the original paper and adopt a mixup coefficient of $\alpha=0.1$. For adversarial training baselines, we directly use the default hyperparameters released by their authors to avoid unintended deviations from the original configurations.
To support reproducibility, we provide the complete training pipeline 4, including data preprocessing scripts, model configuration files (e.g., batch size, sequence length, and optimizer parameters), and execution instructions. This ensures that all experimental results presented in this work can be reproduced under the same computational environment and hyperparameter settings.
6.4 Evaluation metrics and Baselines
@@
In the evaluation, we assess the performance of trained models from two perspectives, namely, clean accuracy and model robustness. For the robustness evaluation, we check the attack success rate (ASR) metric of two well-established adversarial attack methods, namely, ALERT Yang et al. (2022), which uses variable renaming while considering natural semantics, and MHM Zhang et al. (2020), which also randomly replaces variable names but does not account for natural semantics. Metric ASR evaluates the performance of models, including CodeBERT, GraphCodeBERT, and CodeT5 Dong et al. (2025). In addition, we evaluate the natural robustness of all trained models by applying 18 different code refactoring methods to generate diverse program variants. Specifically, following the existing work Dong et al. (2025, 2023), we construct transformed versions of each code sample in the dataset by randomly applying one of the refactoring methods.
@@
We consider model training without data augmentation (No Aug), Refactor Yu et al. (2022), a family of text-oriented data augmentation methods Dong et al. (2025), and MixCode as baseline methods. Refactor, which keeps code syntax rules, is a widely used data augmentation framework specifically designed for code data. In contrast, text-oriented data augmentation methods, which slightly break syntax rules, have recently proven more effective than Refactor. Additionally, MixCode is a more recent and high-performing code augmentation method that significantly enhances code models.
7 Results Analysis
7.1 RQ1: How effective is GenCode in producing accurate code models?
@@
Overall accuracy. Column Accuracy in Table 3 summarizes the final test accuracy of code models trained by using different code augmentation methods including GenCode. Surprisingly, from the results, we can see that most of the existing code augmentation methods except RS and MixCode negatively affect the performance of code models. This finding is similar to the conclusion from recent code augmentation studies Bielik and Vechev (2020); Yu et al. (2022); Dong et al. (2025) which states code augmentation has limited advantage in boosting the training of code models and also indicates that data augmentation for code is a challenging problem.
Table 3: Effectiveness of data augmentation methods w.r.t. test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data. No Aug: without data augmentation. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
Model DA method GCJ Refactory Java250 BigCloneBench
No Aug 90.96 ± 0.07 96.21 ± 0.13 96.36 ± 0.09 96.91 ± 0.12
Refactor 91.74 ± 0.11 95.95 ± 0.19 96.44 ± 0.12 96.97 ± 0.21
SR 70.66 ± 0.11 96.71 ± 0.11 96.21± 0.02 97.02 ± 0.15
RI 59.91 ± 0.21 96.43 ± 0.18 96.15 ± 0.06 97.38 ± 0.23
RS 93.24 ± 0.13 96.73 ± 0.15 96.49 ± 0.04 96.92 ± 0.12
RD 76.98 ± 0.09 96.46 ± 0.18 96.61 ± 0.12 97.11 ± 0.19
BT 82.74 ± 0.15 94.31 ± 0.38 96.06 ± 0.06 97.13 ± 0.17
MixCode 92.21 ± 0.25 96.76 ± 0.31 96.62 ± 0.23 -
CodeBERT GenCode 96.23 ± 0.18 99.32 ± 0.14 98.89 ± 0.09 98.93 ± 0.11
No Aug 93.96 ± 0.15 96.84 ± 0.16 96.41 ± 0.11 96.88 ± 0.09
Refactor 91.72 ± 0.21 95.49 ± 0.21 96.53 ± 0.15 97.11 ± 0.23
SR 89.46 ± 0.13 95.57 ± 0.13 96.46 ± 0.14 97.14 ± 0.16
RI 80.47 ± 0.21 94.51 ± 0.25 96.52 ± 0.09 97.35 ± 0.17
RS 94.76 ± 0.15 97.88 ± 0.15 96.44 ± 0.14 97.54 ± 0.12
RD 91.71 ± 0.26 96.53 ± 0.19 96.56 ± 0.21 97.16 ± 0.21
BT 80.47 ± 0.05 95.99 ± 0.27 96.39 ± 0.13 97.11 ± 0.16
MixCode 94.19 ± 0.25 97.29 ± 0.15 96.89 ± 0.19 -
GraphCodeBERT GenCode 97.93 ± 0.14 99.45 ± 0.17 99.16 ± 0.13 99.23 ± 0.08
No Aug 95.02 ± 0.07 97.22 ± 0.17 96.98 ± 0.08 97.38 ± 0.13
Refactor 92.21 ± 0.14 96.43 ± 0.22 96.92 ± 0.11 97.24 ± 0.12
SR 90.14 ± 0.21 94.67 ± 0.14 96.53 ± 0.05 97.31 ± 0.08
RI 78.77 ± 0.05 95.02 ± 0.15 96.11 ± 0.09 97.46 ± 0.13
RS 95.66 ± 0.14 97.97 ± 0.29 96.34 ± 0.11 97.03 ± 0.15
RD 92.01 ± 0.11 97.52 ± 0.26 97.21 ± 0.07 97.43 ± 0.07
BT 81.45 ± 0.13 95.38 ± 0.11 96.45 ± 0.14 97.65 ± 0.14
MixCode 94.32 ± 0.16 97.32 ± 0.13 96.91 ± 0.11 -
CodeT5 GenCode 98.85 ± 0.11 99.64 ± 0.06 99.38 ± 0.11 99.34 ± 0.12
No Aug 98.91 ± 0.02 98.12 ± 0.01 97.83 ± 0.02 98.75 ± 0.03
Refactor 96.92 ± 0.01 97.99 ± 0.03 97.25 ± 0.02 98.61 ± 0.01
SR 92.36 ± 0.04 97.27 ± 0.04 97.27 ± 0.03 98.46 ± 0.04
RI 90.47 ± 0.05 96.82 ± 0.03 97.26 ± 0.02 98.31 ± 0.02
RS 97.66 ± 0.06 97.41 ± 0.06 98.19 ± 0.01 98.15 ± 0.01
RD 98.33 ± 0.03 97.02 ± 0.02 97.72 ± 0.02 98.36 ± 0.03
BT 89.92 ± 0.02 97.09 ± 0.05 97.41 ± 0.02 98.12 ± 0.05
MixCode 98.56 ± 0.04 98.21 ± 0.08 98.23 ± 0.03 -
StarCoder2-7B GenCode 98.93 ± 0.04 99.65 ± 0.02 99.41 ± 0.05 99.36 ± 0.09
No Aug 98.95 ± 0.03 98.19 ± 0.03 98.11 ± 0.02 98.78 ± 0.01
Refactor 97.22 ± 0.03 97.51 ± 0.02 97.31 ± 0.01 98.32 ± 0.04
SR 92.45 ± 0.02 97.47 ± 0.03 97.37 ± 0.04 98.51 ± 0.05
RI 90.72 ± 0.03 96.31 ± 0.02 98.07 ± 0.05 98.34 ± 0.04
RS 96.89 ± 0.05 97.52 ± 0.05 97.31 ± 0.02 98.02 ± 0.06
RD 96.83 ± 0.01 96.89 ± 0.04 97.52 ± 0.03 98.25 ± 0.04
BT 91.07 ± 0.06 97.15 ± 0.03 98.02 ± 0.02 98.06 ± 0.08
MixCode 98.61 ± 0.03 98.46 ± 0.06 98.53 ± 0.04 -
Qwen2.5-Coder-7B GenCode 98.96 ± 0.04 99.65 ± 0.01 99.42 ± 0.04 99.36 ± 0.05
@@
Then, compared to the existing data augmentation methods, the results demonstrate that GenCode consistently enhances the training and achieves the best results across all the considered datasets and models. Concretely, on average, for pre-trained code models, GenCode generates models with 4.52%, 3.05%, 2.56%, and 2.11% higher accuracy than No Aug in GCJ, Refactory, Java250, and BigCloneBench, respectively. Besides, compared to MixCode, GenCode can achieve 4.1%, 2.34%, and 2.33% better results in GCJ, Refactory, and Java250 datasets, respectively. Moving to LLMs, GenCode improves the performance of StarCoder2 across all tasks by up to 0.93% on average, and achieves a 0.84% average improvement for Qwen2.5-Coder.
Convergence speed analysis. Fig. 4 depicts the average test accuracy at each training epoch of CodeBERT using different data augmentation methods. The results demonstrated that GenCode consistently outperforms other baseline methods during the whole training process. We can conclude that the choice of the number of training epochs will not affect the advantages of our method over other methods.
@@
Figure 4: Convergence speed of CodeBERT using different code augmentation methods in each task.
Visualizing principal component analysis. Furthermore, we explore why models trained by GenCode have higher accuracy than models trained using other methods by analyzing the code representation capability of trained code embeddings. Specifically, 1) first, we extract the code embeddings produced by each trained model using all the test data as code features. 2) Then, since the extracted features are high dimensions and hard to analyze, we reduce the dimensions of code data features using the Principal Component Analysis (PCA) Maćkiewicz and Ratajczak (1993) algorithm. 3) After that, we visualize the reduced 2-dimensional vectors. Normally, the features extracted by a well-trained model are more clearly distributed based on the data labels.
@@
Fig. 5 depicts the visualized features extracted from CodeBERT using the Refactory dataset. We can see that the code embeddings trained by using GenCode are more clearly distributed than embedding extracted by other models, i.e., the buggy code and correct code are greatly different in the figure. These results indicate that GenCode can train good code embeddings and, therefore, produce better accurate code models for code understanding tasks. Besides, we computed the average distance between the data points from different classes, i.e., the average distance between red points and blue points in Fig. 5. The results from Table 4 confirm that GenCode can better distinguish the two classes. Interestingly, we found that MixCode has the smallest distance result. That means even though MixCode can train code models with good performance, the code models have low confidence in their predictions, thus, the data are near the decision boundaries. This phenomenon is similar to the findings in Verma et al. (2019), where interpolating hidden state representations yields two beneficial properties: 1) class representations are flattened into a minimal number of directions of variation, and 2) all points between these flat representations are assigned low-confidence predictions.